# Enhancing Diffusion Model Stability for Image Restoration via Gradient Management

### v1

@@ Abstract.
Diffusion models have shown remarkable promise for image restoration by leveraging powerful priors. Prominent methods typically frame the restoration problem within a Bayesian inference framework, which iteratively combines a denoising step with a likelihood guidance step. However, the interactions between these two components in the generation process remain underexplored. In this paper, we analyze the underlying gradient dynamics of these components and identify significant instabilities. Specifically, we demonstrate conflicts between the prior and likelihood gradient directions, alongside temporal fluctuations in the likelihood gradient itself. We show that these instabilities disrupt the generative process and compromise restoration performance. To address these issues, we propose Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique. SPGD integrates two synergistic components: (1) a progressive likelihood warm-up strategy to mitigate gradient conflicts; and (2) adaptive directional momentum (ADM) smoothing to reduce fluctuations in the likelihood gradient. Extensive experiments across diverse restoration tasks demonstrate that SPGD significantly enhances generation stability, leading to state-of-the-art performance in quantitative metrics and visually superior results. Code is available at here.
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelanddoi: XXXXXXX.XXXXXXXsubmissionid: 7272ccs: Computing methodologies Computer vision problemsccs: Computing methodologies Machine learning
@@ 4.1. Progressive Warm-Up Phase
Specifically, within a single conceptual reverse step from time $t$ to $t-1$, starting from the current estimate $\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}$, we perform $N$ iterative refinement steps focusing solely on the likelihood gradient using a reduced step size:
@@ 4.3. Theoretical Analysis
(14) $L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
@@
Table 1. Quantitative evaluation of image restoration tasks on FFHQ 256$\times$256 and ImageNet 256$\times$256 with $\sigma_{y}=0.01$.Best results are marked in bold, and second-best results are underlined.
@@ 5.1. Experimental Setup
Datasets & Metrics. We evaluate our proposed SPGD method and baseline approaches on 1,000 images from the FFHQ 256$\times$256 validation set (Karras et al., 2019) and 500 images from the ImageNet 256$\times$256 validation set (Russakovsky et al., 2015). The pre-trained diffusion models for FFHQ and ImageNet are obtained from (Chung et al., 2023) and (Dhariwal and Nichol, 2021), respectively, following established conventions.
@@
Tasks. We consider four standard linear image restoration tasks: (1) Inpainting: Randomly mask 80% of the pixels across all RGB channels. (2) Gaussian Blur: Apply a Gaussian blur kernel of size $61\times 61$ with a standard deviation of $\sigma=3.0$. (3) Motion Blur: Use a motion blur kernel of size $61\times 61$ and intensity 0.5, following the procedure of Chung et al. (Chung et al., 2022). (4) Super-Resolution (SR) $\times$4: Perform 4×4$\times$4 × bicubic downsampling.
@@ 5.2. Quantitative Results
On the more challenging ImageNet dataset, our method continues to outperform the baselines in most scenarios. It achieves the best SSIM and LPIPS scores for inpainting, Gaussian deblurring, and motion deblurring. For SR×$\times$4, our method obtains the best PSNR and LPIPS, and competitive SSIM scores, demonstrating its effectiveness on diverse image content. While DDRM achieves slightly higher SSIM for SR ($\times$4), the reliance on the singular value decomposition (SVD) of operator $\mathcal{A}$ renders their methodology inapplicable in many scenarios. Our method shows a marked advantage in LPIPS and frequently outperforms others in distortion metrics, demonstrating that our gradient management directly leads to the enhanced stability with superior restoration quality.
@@ 5.4.1. Impact of Design Components
The results, presented in Table 2, reveal the following observations: (i) Implementing only the warm-up strategy yields significant improvements across all metrics compared to the baseline. This confirms that performing multiple, smaller likelihood updates before the denoising step effectively mitigates gradient conflicts and enhances reconstruction quality. (ii) Applying ADM smoothing directly fails to improve performance. We hypothesize this occurs because the momentum accumulates information based on likelihood gradients calculated at different states ($t-1$, $t$, etc.) where the underlying image estimate has significantly changed due to the intervening denoising steps. This makes the historical gradient information less relevant or potentially misleading for the current step’s likelihood correction. (iii) Combining both techniques results in the best performance, substantially outperforming either component alone. This demonstrates a crucial synergy: ADM effectively stabilizes the likelihood gradient within the warm-up phase, where likelihood gradients $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$ are computed at the same effective timestep $t$. These findings underscore that both components are essential and work synergistically, validating the design of SPGD.
5.4.2. Ablation Study on $N$
@@
Table 3. Ablation study on the impact of warm-up steps $N$.
@@ B.1. Descent Lemma for L-Smooth Functions
(22) $f(\mathbf{w}^{\prime})\leq f(\mathbf{w})+\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\frac{L}{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}^{2}$
@@
[…]
@@
(24) $-\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle\,dt$
@@
(25) $\leq$ $\|\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\nabla f(\mathbf{w})\|_{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}$
@@
(26) $=$ $Lt\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}$
@@
[…]
@@
$\sum_{j=0}^{N-1}[L_{t}(\mathbf{x}_{t}^{(j+1)})-L_{t}(\mathbf{x}_{t}^{(j)})]\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
@@
$L_{t}(\mathbf{x}_{t}^{(N)})-L_{t}(\mathbf{x}_{t}^{(0)})\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
@@
$L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
@@ D.2. Our SPGD:
Super-Resolution (×$\times 4$) 8 7
@@ D.3. Computational Cost
Table 6. Computational Cost ($T=100$) under different $N$.
Metric $N$=1 $N$=2 $N$=5 $N$=10 $N$=20

### v2

@@ Abstract.
Diffusion models have shown remarkable promise for image restoration by leveraging powerful priors. Prominent methods typically frame the restoration problem within a Bayesian inference framework, which iteratively combines a denoising step with a likelihood guidance step. However, the interactions between these two components in the generation process remain underexplored. In this paper, we analyze the underlying gradient dynamics of these components and identify significant instabilities. Specifically, we demonstrate conflicts between the prior and likelihood gradient directions, alongside temporal fluctuations in the likelihood gradient itself. We show that these instabilities disrupt the generative process and compromise restoration performance. To address these issues, we propose Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique. SPGD integrates two synergistic components: (1) a progressive likelihood warm-up strategy to mitigate gradient conflicts; and (2) adaptive directional momentum (ADM) smoothing to reduce fluctuations in the likelihood gradient. Extensive experiments across diverse restoration tasks demonstrate that SPGD significantly enhances generation stability, leading to state-of-the-art performance in quantitative metrics and visually superior results. Code is available at https://github.com/74587887/SPGD.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754582isbn: 979-8-4007-2035-2/2025/10submissionid: 7272ccs: Computing methodologies Computer vision problemsccs: Computing methodologies Machine learning
@@ 4.1. Progressive Warm-Up Phase
Specifically, within a single conceptual reverse step from time $t$ to t1t-1, starting from the current estimate $\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}$, we perform $N$ iterative refinement steps focusing solely on the likelihood gradient using a reduced step size:
@@ 4.3. Theoretical Analysis
(14) $L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2},$
@@
Table 1. Quantitative evaluation of image restoration tasks on FFHQ $\times256$ and ImageNet $\times256$ with $\sigma_{y}=0.01$.Best results are marked in bold, and second-best results are underlined.
@@ 5.1. Experimental Setup
Datasets & Metrics. We evaluate our proposed SPGD method and baseline approaches on 1,000 images from the FFHQ $\times256$ validation set (Karras et al., 2019) and 500 images from the ImageNet $\times256$ validation set (Russakovsky et al., 2015). The pre-trained diffusion models for FFHQ and ImageNet are obtained from (Chung et al., 2023) and (Dhariwal and Nichol, 2021), respectively, following established conventions.
@@
Tasks. We consider four standard linear image restoration tasks: (1) Inpainting: Randomly mask 80% of the pixels across all RGB channels. (2) Gaussian Blur: Apply a Gaussian blur kernel of size $61\times 61$ with a standard deviation of $\sigma=3.0$. (3) Motion Blur: Use a motion blur kernel of size $61\times 61$ and intensity 0.5, following the procedure of Chung et al. (Chung et al., 2022). (4) Super-Resolution (SR) $\times$4: Perform $4\times$ bicubic downsampling.
@@ 5.2. Quantitative Results
[…]
@@ 5.4.1. Impact of Design Components
The results, presented in Table 2, reveal the following observations: (i) Implementing only the warm-up strategy yields significant improvements across all metrics compared to the baseline. This confirms that performing multiple, smaller likelihood updates before the denoising step effectively mitigates gradient conflicts and enhances reconstruction quality. (ii) Applying ADM smoothing directly fails to improve performance. We hypothesize this occurs because the momentum accumulates information based on likelihood gradients calculated at different states (t1t-1, $t$, etc.) where the underlying image estimate has significantly changed due to the intervening denoising steps. This makes the historical gradient information less relevant or potentially misleading for the current step’s likelihood correction. (iii) Combining both techniques results in the best performance, substantially outperforming either component alone. This demonstrates a crucial synergy: ADM effectively stabilizes the likelihood gradient within the warm-up phase, where likelihood gradients $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$ are computed at the same effective timestep $t$. These findings underscore that both components are essential and work synergistically, validating the design of SPGD.
5.4.2. Ablation Study on NN
@@
Table 3. Ablation study on the impact of warm-up steps NN.
@@ B.1. Descent Lemma for L-Smooth Functions
(22) $f(\mathbf{w}^{\prime})\leq f(\mathbf{w})+\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\frac{L}{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}^{2}.$
@@
(23) $f(\mathbf{w}^{\prime})-f(\mathbf{w})=\int_{0}^{1}\langle\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w})),\mathbf{w}^{\prime}-\mathbf{w}\rangle\,dt.$
Adding and subtracting $\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle=\int_{0}^{1}\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle\,dt$:
@@
(24) $-\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle\,dt.$
@@
(25) $\leq$ $\|\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\nabla f(\mathbf{w})\|_{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}.$
@@
(26) $=$ $Lt\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}.$
@@
$\leq$ $\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\int_{0}^{1}(Lt\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2})\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}\,dt$
(27) $=$ $\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+L\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}^{2}\int_{0}^{1}t\,dt.$
Evaluating the integral $\int_{0}^{1}t\,dt=1/2$, we obtain:
@@
$\sum_{j=0}^{N-1}[L_{t}(\mathbf{x}_{t}^{(j+1)})-L_{t}(\mathbf{x}_{t}^{(j)})]\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}.$
@@
$L_{t}(\mathbf{x}_{t}^{(N)})-L_{t}(\mathbf{x}_{t}^{(0)})\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}.$
@@
$L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}.$
@@ D.2. Our SPGD:
Super-Resolution ($\times 4$) 8 7
@@ D.3. Computational Cost
Table 6. Computational Cost ($T=100$) under different NN.
Metric NN=1 NN=2 NN=5 NN=10 NN=20